log(sqrt(x),27)=-6
Answer by nerdybill(7384)

log(sqrt(x),27)=-6
I'm assuming you mean (on the left), log base sqrt(x) of 27
.
log(27)/log(sqrt(x)) = -6
log(27) = -6*log(sqrt(x))
-log(27)/6 = log(sqrt(x))
-0.2386 = log(sqrt(x))
10^(-0.2386) = sqrt(x)
10^(-0.2386) = sqrt(x)
0.5774 = sqrt(x)
0.3333 = x
or
1/3 = x